OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (6, −8) and has a slope of −2? @phi

OpenStudy (anonymous):

@Destinymasha

OpenStudy (anonymous):

@Opcode

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

@Donblue22 plz help when possible

OpenStudy (anonymous):

@TheSmartOne help PLEASE! ;-;

TheSmartOne (thesmartone):

Point slope form is y=mx+b

TheSmartOne (thesmartone):

We know the slope is -2 So it becomes y=-2x+b

TheSmartOne (thesmartone):

We know that the point (6,-8) works in it so lets plug it in to find b. -8=-2(6)+b -8=-12+b b=4

TheSmartOne (thesmartone):

So the final equation is y=-2x+4

TheSmartOne (thesmartone):

@thundernova308

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@TheSmartOne Just looked and is not a naswer choice ;.;

TheSmartOne (thesmartone):

What are your answer choices. Did you type your question correctly?